OpenStudy (anonymous):

Which number is a common factor of 36, 48, and 66? (Points : 1) 12 10 8 6

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Hint: How do you find the factors of a number?

OpenStudy (skullpatrol):

Look for the number that divides into all three evenly.

OpenStudy (dahlindubs11):

does 12 go into all three numbers? 10? 8? 6? which of the choices go into the three numbers?

OpenStudy (welshfella):

you can also do this by writing them in prime factor from and looking for common factors 36 = 2 * 2 * 3 * 3 48 = 2 * 2 * 2* 2 * 3 66 = 2*3* 11 notice that is one triplicate of 2 and one triplicate of 3 so GCF = 2*3 = 6